# Reference.bib

### v1

@@ 5.2. Connected Limiting Fiber (CLF)
$A\subset Z_{\mathfrak{C}(J)}$.
Proposition 5.9 will follow from the construction of $\mathfrak{C}(J)$ and the following lemma:
Lemma 5.10.
If $\overline{A_{J}}\cap Z_{K}\neq\emptyset$, then $\boldsymbol{\tau}_{K}\leq\boldsymbol{\tau}_{J}$. Moreover if $\overline{A_{J}}\cap\overline{A_{K}}\neq\emptyset$, then $J\sim K$, $J\sim J\cup K$ and $K\sim J\cup K$.
@@
Near any $p\in\overline{A_{J}}\cap\overline{A_{K}}\subset Z_{J\cup K}$, choose a local coordinate neighborhood $B_{p}$ centered at $p$ admitting diagram (3.5) with local coordinates $\{\vec{\xi}=\{\xi_{1},...,\xi_{r}\},\vec{\eta}=\{\eta_{1},...,\eta_{s}\},\vec{\mu}=\{\mu_{1},...,\mu_{m}\}\}$ such that
$Z_{J}\cap B_{p}=\{\mu_{1}=...=\mu_{m}=0\},$
$Z_{J\cup K}\cap B_{p}=\{\eta_{1}=...=\eta_{s}=\mu_{1}=...=\mu_{m}=0\}.$
By (3.5), over $A\cap Z_{J}^{*}$ the local lift of the boundary period map has the form:
$\widetilde{\Phi_{p}}:(\vec{\xi},\vec{\eta},0)\rightarrow\exp(\sum_{j\in J}\frac{\log(\xi_{j})}{2\pi i}N_{j}+\sum_{k\in K-J}\frac{\log(\eta_{k})}{2\pi i}N_{k})\psi_{p}(\xi,\eta,0).$
For some holomorphic map $\psi_{p}:B_{p}\rightarrow\check{D}$.
Since $A_{J}$ is a component of some CLF $A$, the image of $\Phi_{p}|_{A_{J}}$ takes value in a single $\exp(\mathbb{C}\tau_{J})$-orbit in $\check{D}$, say it is $\exp(\mathbb{C}\tau_{J})F_{0}$. Since $p\in\overline{A_{J}}\cap Z_{J\cup K}$, there exist sequences $\{\vec{\xi^{l}}\},\{\vec{\eta^{l}}\}\rightarrow\vec{0}$ such that
$\widetilde{\Phi_{p}}(\vec{\xi}^{l},\vec{\eta}^{l},0)=\exp(\sum_{j\in J}\frac{\log(\xi^{l}_{j})}{2\pi i}N_{j}+\sum_{k\in K-J}\frac{\log(\eta^{l}_{k})}{2\pi i}N_{k})\psi_{p}(\xi^{l},\eta^{l},0)\subset\exp(\mathbb{C}\tau_{J})F_{0}.$
This implies $\psi_{p}(\xi^{l},\eta^{l},0)\in\exp(\mathbb{C}\tau_{J})F_{0}$ and $\sum_{j\in J}\frac{\log(\xi^{l}_{j})}{2\pi i}N_{j}+\sum_{k\in K-J}\frac{\log(\eta^{l}_{k})}{2\pi i}N_{k}$ always takes value in $\mathbb{C}\tau_{J}$ regardless of the choice of branches of logarithm. Since $\{N_{j},j\in J\}\subset\sigma_{J}\subset\tau_{J}$, we must have $\{N_{k},k\in K-J\}\subset\tau_{J}$, which implies $\tau_{J\cup K}=\tau_{J}$ hence $J\sim J\cup K$ and $\tau_{K}\leq\tau_{J}$. This proves the first part. For the second part we also have $K\sim J\cup K$ and $J\sim K$. ∎
5.3. Monodromy around a CLF
For every $p\in A$ we choose the neighborhood $B_{p}$ as in the diagram (5.2). Since $A$ is compact, we may choose a finite open cover $B_{A}:=\bigcup_{k=1,...,n}B_{p_{k}}$ of $A$, and let $\mathcal{U}_{A}^{-}:=B_{A}\cap(B\cup Z_{\mathfrak{C}(J)})$. Denote $\Gamma_{A}$ as the monodromy group of $\mathbb{V}\rightarrow\mathcal{U}_{A}$.
Lemma 5.11.
$\Gamma_{A}=\Gamma_{\tau_{J}}$, in particular $\Gamma_{A}$ is abelian.
Proof.
The idea is similar to [GGR21, Lemma 4.15]. Since $A$ is a CLF, after choosing proper lifts of local period maps in (5.2), we may find a unique nilpotent orbit $(\tau_{J},F)$ such that for each $p_{k}\in A$, the local KNU map $\Phi_{p_{k}}^{\mathrm{KNU}}$ on $\mathcal{U}_{p_{k}}^{-}$ sends $p_{k}$ to $(\tau_{J},F)$. Different local lifts of $\Phi_{p_{k}}^{\mathrm{KNU}}$ satisfying this condition are conjugated by the stabilizer of $(\tau_{J},F)$ which is exactly $\Gamma_{\tau_{J}}$ according to [KU08, Prop. 7.4.3]. This implies we have a well-defined map
$\Phi_{A}:\mathcal{U}_{A}\rightarrow\Gamma_{\tau_{J}}\backslash D$
which implies $\Gamma_{A}\leq\Gamma_{\tau_{J}}$. The other direction of inclusion is evident. ∎
Corollary 5.12.
The local extended period maps $\Phi_{p}^{\mathrm{KNU}}$ for every $p\in A$ can be patched together and extended to a map
(5.3) $\Phi_{A}^{\mathrm{KNU}}:B_{A}\rightarrow\Gamma_{\tau_{J}}\backslash D_{\tau_{J}}$
whose restriction on $\mathcal{U}_{A}$ is the original period map.
Proof.
By patching $\Phi_{p}^{\mathrm{KNU}}$ in (5.2) for every $p\in A$ we get a well-defined map $\Phi_{A}^{\mathrm{KNU}}:\mathcal{U}_{A}^{-}\rightarrow\Gamma_{\tau_{J}}\backslash D_{\tau_{J}}$. Shrink $B_{A}$ if necessary, we may assume $B_{A}\cap Z_{K}\neq\emptyset$ if and only if $Z_{K}\cap A\neq\emptyset$. In this case, by Lemma 5.10 we have $\boldsymbol{\tau}_{K}\leq\boldsymbol{\tau}_{J}$ and we may extend $\Phi_{A}^{\mathrm{KNU}}$ from $\mathcal{U}_{A}^{-}$ to $B_{A}$. ∎
5.4. Proof of Theorem 5.1
The first step is to construct the cover $\overline{B}=\bigcup_{k}\hat{U}_{k}$. Let $A\subset\overline{B}$ be a CLF. By Corollary 5.12 there is a neighborhood $B_{A}\subset\overline{B}$ over which there exists a KNU map $\Phi_{A}^{\mathrm{KNU}}$. Moreover, $A\subset B_{A}$ is the $\Phi_{A}^{\mathrm{KNU}}$-fiber of some $(\tau_{J},F_{A})\in\Gamma_{\tau_{J}}\backslash D_{\tau_{J}}$. We may choose a neighborhood $(\tau_{J},F_{A})\in V_{A}\subset\Gamma_{\tau_{J}}\backslash D_{\tau_{J}}$ such that $(\Phi_{A}^{\mathrm{KNU}})^{-1}(V_{A})=:W_{A}\subset\overline{W_{A}}\subset B_{A}$. Such $W_{A}$ must exist as $\overline{B}$ is a $T_{3}$-topological space. By the description of $\Gamma_{\tau_{J}}\backslash D_{\tau_{J}}$ as a logarithmic manifold [KU08, Chap. 3], we may further shrink $V_{A}$ and corresponding $W_{A}$ such that there exists a complex analytic space $V_{A}^{\prime}$ as well as a morphism of fs logarithmic analytic spaces $V_{A}\hookrightarrow V_{A}^{\prime}$ such that the composition map $W_{A}\rightarrow V_{A}^{\prime}$ is a morphism of complex spaces.
Take the cover of $\overline{B}$ by open sets $W_{A}$ as $A$ running over every CLF, and take a finite open subcover $\overline{B}=\bigcup_{1\leq k\leq n}W_{k}$ with respect to CLF’s $\{A_{1},...,A_{n}\}$.
Claim 5.13.
$\overline{B}=\bigcup_{1\leq k\leq n}W_{k}$ satisfies all properties required by Theorem 5.1.
Over each $W_{k}$, we denote the map given by Corollary 5.12 as
$\Phi_{k}^{\mathrm{KNU}}:W_{k}\rightarrow\Gamma_{k}\backslash D_{\sigma_{k}}.$
It remains to show $W_{k}$ and the maps $\Phi_{k}^{\mathrm{KNU}}$ satisfy the required properties.
For conditions (1) and (2) of Theorem 5.1, thanks to Lemma 5.11 and Corollary 5.12, the only remaining thing to be proved is the properness of maps $\Phi_{k}^{\mathrm{KNU}}$.
Denote $q_{k}:=\Phi_{k}^{\mathrm{KNU}}(A_{k})\in\Gamma_{k}\backslash D_{\sigma_{k}}$. By the choice of $W_{k}$, there exists complex analytic spaces $V_{k}^{\prime}$ such that each $\Phi_{k}^{\mathrm{KNU}}$ may be regarded as a morphism of complex spaces:
$\Phi_{k}^{\mathrm{KNU}}:W_{k}\rightarrow V_{k}\hookrightarrow V_{k}^{\prime}.$
We may take a closed neighborhood $q\in F_{k}\subset V_{k}^{\prime}$ such that $(\Phi_{k}^{\mathrm{KNU}})^{-1}(F_{k})$ is closed in $W_{k}$. This implies up to shrinking $W_{k}$ to $(\Phi_{k}^{\mathrm{KNU}})^{-1}(F_{k}^{\circ})$, we may assume $\Phi_{k}^{\mathrm{KNU}}$ is proper.
Condition (3) directly follows from the definition. We now show (4). It is enough to show that for any smooth point $p\in\mathcal{F}_{k}\cap\mathcal{F}_{l}\cap Z_{K}^{*}$, $\mathcal{F}_{k}$ and $\mathcal{F}_{l}$ agree in some small neighborhood $\Delta_{p}$ of $p$. This is immediate, as if we choose $\Delta_{p}\subset Z_{K}^{*}$ to be contractible, both $\mathcal{F}_{k}\cap\Delta_{p}$ and $\mathcal{F}_{l}\cap\Delta_{p}$ are just the $\Phi_{p}^{\mathrm{KNU}}$-fiber in $\Delta_{p}$ containing $p$. The proof of Theorem 5.1 is concluded.
6. Construction of the analytic structure
In this section, we will finish the construction purposed by Theorem 2.8 based on Theorem 5.1.
Assume all ingredients provided by Theorem 5.1 are fixed. These include a finite cover $\overline{B}=\bigcup_{k}W_{k}$, and for every $k$, a local KNU extension $\overline{\Phi}_{k}:W_{k}\rightarrow V_{k}\hookrightarrow\Gamma_{k}\backslash D_{\tau_{k}}$ which is proper. Theorem 3.12 implies that the image of each $\overline{\Phi}_{k}$ is complex analytic. We may assume every $\overline{\Phi}_{k}:W_{k}\rightarrow V_{k}$ has connected fibers, this can always be achieved after applying a Stein factorization.
On each $k$, there is a complex equivalence relation on $W_{k}$ defined by $W_{k}\times_{V_{k}}W_{k}\subset W_{k}\times W_{k}$. Since each $\overline{\Phi}_{k}$ is a proper holomorphic map, the equivalence relation $W_{k}\times_{V_{k}}W_{k}$ is proper holomorphic in the sense of [Gra83].
Note that $\bigcup_{k}(W_{k}\times W_{k})\subset\overline{B}\times\overline{B}$ is an open analytic subspace containing the diagonal.
Proposition 6.1.
$\bigcup_{k}(W_{k}\times_{V_{k}}W_{k})\subset\bigcup_{k}(W_{k}\times W_{k})\subset\overline{B}\times\overline{B}$ gives a proper holomorphic equivalence relation on $\overline{B}$ with connected fibers whose restriction on each $(W_{k}\times W_{k})$ agrees with $W_{k}\times_{V_{k}}W_{k}$.
Proof.
For every $k$ let $E_{k}^{\prime}:=W_{k}\times W_{k}$ and $E_{k}:=W_{k}\times_{V_{k}}W_{k}$. We first show that for every $k,l$, $E_{k}^{\prime}\cap E_{l}=E_{l}^{\prime}\cap E_{k}$. Since for any $(a,b)\in E_{k}^{\prime}$, $(a,b)\in E_{k}$ if and only if $a,b$ lie in the same fiber of $\overline{\Phi}_{k}$, it suffices to show that on $W_{k}\cap W_{l}$, the fibers of $\overline{\Phi}_{k}$ and $\Phi_{l}^{\mathrm{KNU}}$ agree. This is exactly (2) and (4) of Theorem 5.1 by our assumption that all $\overline{\Phi}_{k}$ have connected fibers. The proper and holomorphic conditions follow from the fact that both conditions are locally around a fiber of $\bigcup_{k}(W_{k}\times_{V_{k}}W_{k})\subset\overline{B}$, which is consistent with the fibers for each $E_{k}\subset E_{k}^{\prime}$. ∎
Denote $E\subset\overline{B}\times\overline{B}$ as the resulting proper holomorphic equivalence relation by Proposition 6.1, and let ¯$\overline{\wp}:=\overline{B}/\sim_{E}$ as the quotient space and let $q:\overline{B}\rightarrow\overline{\wp}$ be the quotient map. By [Gra83, Theorem 2], $\overline{\wp}$ admits a structure of complex analytic space. More precisely, the structure sheaf $\mathcal{O}_{\overline{\wp}}$ of $\overline{\wp}$ is given by pushing forward the analytic germs of $\mathcal{O}_{\overline{B}}$ constant on every $\overline{\Phi}_{k}$-fiber.
Lemma 6.2.
$E^{*}:=E\cap(B\times B)$ gives the equivalence relation on $B$ defining the (Stein factorization of) period map $\Phi^{\prime}:B\rightarrow\wp^{\prime}$.
Proof.
It is enough to show $\Phi^{\prime}$ has the same fibers as $E$. Take a fiber $\zeta$ of $\Phi^{\prime}$, then $\zeta\subset B$ is connected, compact and hence closed in $\overline{B}$. For any $k$ such that $U_{k}\cap\zeta\neq\emptyset$, since the period map $\Phi_{k}:U_{k}\rightarrow\Gamma_{k}\backslash D$ is a lift of $\Phi:B\rightarrow\Gamma\backslash D$ when restricted to $U_{k}$, $\zeta_{k}:=\zeta\cap U_{k}$ is contained in a unique fiber of $\Phi_{k}$. But by the construction of Theorem 5.1, the fibers of $\Phi_{k}$ are compact and contained in $U_{k}$, this means $\zeta_{k}\cap\partial U_{k}=\emptyset$, hence we must have $\zeta\subset U_{k}$. This implies $\zeta$ is a fiber of $\Phi^{\prime}$ if and only if it is a fiber of some $\Phi_{k}$ and hence a fiber of $E$. ∎
Lemma 6.3.
Let $\zeta\subset Z_{I}^{*}\subset\overline{B}$ be a connected analytic subvariety. Then $\zeta$ is contained in a fiber of $E$ if and only if for all $b\in\zeta$ the nilpotent orbit class of $(\tau_{I},F_{b})$ given by Corollary 3.4 and the CMCX $\mathcal{T}$ is constant.
Proof.
If $\zeta$ is contained in a fiber of $E$, it is also contained in the fiber of some $\overline{\Phi}_{k}:W_{k}\rightarrow V_{k}\hookrightarrow\Gamma_{k}\backslash D_{\tau_{k}}$. On the other hand, if over $\zeta$ the nilpotent orbit class $(\tau_{I},F_{b})$ is constant, $\zeta\cap W_{k}$ is contained in a $\overline{\Phi}_{k}$-fiber for any $k$, hence $\overline{\zeta}$ must be contained in every $W_{k}$ with $W_{k}\cap\zeta\neq\emptyset$. In this case $\zeta$ is contained in a fiber of $\overline{\Phi}_{k}$ and hence a fiber of $E$. ∎
To sum up, the space $\overline{\wp}$ and the quotient map $q:\overline{B}\rightarrow\overline{\wp}$ satisfy the requirement of Theorem 2.8, thus $q=\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ is the purposed completed period map of the Stein factorization $B\xrightarrow{\Phi^{\prime}}\wp^{\prime}\rightarrow\wp\subset\Gamma\backslash D$.
7. Examples of projective completions
Let us assume $B$ is a quasi-projective surface. The $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ gives the period map $\Phi:B\rightarrow\Gamma\backslash D$. In this section, we do not assume the monodromy group $\Gamma$ is neat as this is the most common case in practice. Moreover, we fix a projective completion $\overline{B}$ of $B$ which we do not assume to be a SNC completion. Let $B^{\circ}\subset\overline{B}$ be the complement of codimension-$2$ strata. The main result of this section is:
Proposition 7.1.
If there is no CLF (Definition 5.8) of positive dimension in $B^{\circ}$ exist, then we may choose the completed period map $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ constructed by Theorem 2.8 to lie in the category of projective varieties.
Proof.
Let $\varphi:B^{\prime}\rightarrow B$ be a finite etále morphism such that the lift $\mathbb{V}^{\prime}\rightarrow B^{\prime}$ of $\mathbb{V}\rightarrow B$ has neat monodromy group $\Gamma^{\prime}\leq\Gamma$. We may find a projective completion $\overline{\varphi}:\overline{B^{\prime}}\rightarrow\overline{B}$ of $\varphi$ which is a finite morphism. This can be done by normalizing $\overline{B}$ in the function field of $B^{\prime}$. Blowing up $\overline{B^{\prime}}$ to some $\overline{B}$ such that $\overline{B}$ is an SNC completion of $B^{\prime}$. Through the composition $\overline{B}\rightarrow\overline{B^{\prime}}\rightarrow\overline{B}$, any positive-dimensional CLF in $\overline{B}$ must be mapped to a codimensional-$2$ point in $\overline{B^{\prime}}$ according to the assumption. Therefore, the completed period map $\overline{\Phi}:\overline{B}\rightarrow\overline{\wp}$ is described by contracting a (possibly empty) set of exceptional divisors of $\overline{B}\rightarrow\overline{B^{\prime}}$. As a consequence, $\overline{\wp}$ acquires its projective structure from some intermediate variety of $\overline{B}\rightarrow\overline{B^{\prime}}$. ∎
This criteria is powerful in practice: It avoids the steps finding a lift with neat monodromy as well as an SNC completion of $B$.
@@ 7.1. Hosono–Takagi’s example
[…]
@@
By blowing up $\mathbb{P}^{2}$ at each tangent point for $5$ times, we get a smooth projective completion $\overline{B}$ of $B$ with SNC boundary divisors. We mark all local monodromy nilpotent cones as follows:
@@
[…]
@@
In $\overline{B}$ the only possible positive-dimensional CLF are the exceptional divisors of $\overline{B}\rightarrow\mathbb{P}^{2}$.
@@
There are no curves in $B$ contracted by $\Phi_{0}$ as in $\mathbb{P}^{2}$ the closure of any such curve must intersect the boundary divisor $\mathrm{Dis}$ while no points on $\mathrm{Dis}$ give the trivial monodromy. The fact that nilpotent cones of type (1) and (3) listed above are of two-dimensional implies non of $D_{0},D_{1},D_{2}$ or $C$ can be a $1$-dimensional CLF. ∎
As a consequence, if we choose $\mathbb{P}^{2}$ instead of $\overline{B}$ to be the projective completion of $B$, the assumption of Proposition 7.1 is met and we may conclude the period map $\Phi_{0}$ (after being lifted to a finite etále cover) admits a projective completion by Theorem 2.8.
@@ 7.2. More examples
In the paper [CD24], Chongyao Chen and the author have investigated several $2$-parameter families of Calabi-Yau threefolds defined over normal rational surfaces as well as their period mappings. The LMHS types along each boundary strata are analyzed in detailed in [CD24, Sec. 4]. We claim that all of these families satisfy the criteria given by Proposition 7.1, therefore the corresponding period mappings admit projective completions. We left the details to readers.
@@ 8. Algebraicity of weakly special subvarieties
In this section we introduce another application of our main theorem. It is well-known that weakly special subvarieties associated to the period mapping $\Phi:B\rightarrow\Gamma\backslash D$ are algebraic subvarieties of $B$. Its proof uses the definability of period mappings in o-minimal geometry and functional transcendence [BKT20], [BBT22]. We will give a new proof of this result without using o-minimal geometry.
Regarding the necessary backgrounds, we refer readers to [KO21] or [Kl22] for a comprehensive overview.

### v4

@@ 6. Examples of projective completions in dimension 22
Let $\overline{S}\to\hat{S}$ be a birational morphism, such that $\overline{S}$ is a smooth SNC completion of $S$, and $S^{\circ}\subset\overline{S}$. Assume that every CLF in $\overline{S}$ maps to a point in $\hat{S}$. Then the the completed period map $\overline{\Phi}:\overline{S}\rightarrow\overline{\wp}$ constructed by Theorem 2.8 has the property that $\overline{\wp}$ is projective.
@@
Through the blow-down map $\pi:\overline{S}\rightarrow\hat{S}$, any positive-dimensional CLF in $\overline{S}$ must be mapped to a codimensional-$2$ point in $\hat{S}$ according to the assumption. Therefore, since the completed period map $\overline{\Phi}:\overline{S}\rightarrow\overline{\wp}$ has connected fibers, we see that $\pi$ factors as $\overline{S}\to\overline{\wp}\to\hat{S}$. The claim now follow by [Abr+02, Thm 2.6.2.].
Proposition 6.1 will be used to justify the projectivity of the extended period map in the following example.
@@ 6.1. Hosono–Takagi’s example
We look at the family of Calabi-Yau threefolds studied by [HT14], [HT18], [Den22] and [DR23]. The period map used here is the one introduced by [HT14] and [HT18], which is a period map of type $(1,2,2,1)$ as follows:
@@
Let $\varphi:S^{\prime}\rightarrow S$ be a finite etále morphism such that the lift $\mathbb{V}^{\prime}\rightarrow S^{\prime}$ of $\mathbb{V}\rightarrow S$ has neat monodromy group $\Gamma^{\prime}\leq\Gamma$. We may find a projective completion $\overline{\varphi}:\overline{S^{\prime}}\rightarrow\mathbb{P}^{2}$ of $\varphi$ which is a finite morphism. This can be done by normalizing $\mathbb{P}^{2}$ in the function field of $S^{\prime}$. As a consequence, the $\mathbb{Z}$-PVHS $\mathbb{V}^{\prime}\rightarrow S^{\prime}$ and $S^{\prime}\subset\overline{S^{\prime}}$ satisfy the condition in Proposition 6.1 if and only if $\mathbb{V}\rightarrow S$ and $S\subset\mathbb{P}^{2}$ do.
By blowing up $\mathbb{P}^{2}$ at each tangent point for $5$ times, we get a smooth projective completion $\overline{S}$ of $S$ with SNC boundary divisors. We mark all local monodromy nilpotent cones as follows:
@@
    $\tau_{j},\ 1\leq j\leq 6$ are all of the type $\langle\mathrm{I}_{1}|\mathrm{I}_{2}|\mathrm{I}_{1}\rangle$.
@@
In $\overline{S}$ the only possible positive-dimensional CLF are the exceptional divisors of $\overline{S}\rightarrow\mathbb{P}^{2}$.
@@
There are no curves in $S$ contracted by $\Phi_{0}$ as in $\mathbb{P}^{2}$ the closure of any such curve must intersect the boundary divisor $\mathrm{Dis}$ while no points on $\mathrm{Dis}$ give the trivial monodromy. The fact that nilpotent cones of type (1) and (3) listed above are of two-dimensional implies non of $D_{0},D_{1},D_{2}$ or $C$ can be a $1$-dimensional CLF. ∎
As a consequence, if we choose $\mathbb{P}^{2}$ instead of $\overline{S}$ to be the projective completion of $S$, the assumption of Proposition 6.1 is met and we may conclude the period map $\Phi_{0}$ (after being lifted to the finite etále cover $S^{\prime}\rightarrow S$) admits a projective completion by Theorem 2.8.
@@ 6.2. More examples
In the paper [CD24], Chongyao Chen and the first author have investigated several $2$-parameter families of Calabi-Yau threefolds defined over normal rational surfaces as well as their period mappings. The LMHS types along each boundary strata are analyzed in detailed in [CD24, Sec. 4]. We claim that all of these families satisfy the criteria given by Proposition 6.1, therefore the corresponding period mappings admit projective completions. We left the details to readers.
@@ 7. Algebraicity of weakly special subvarieties
In this section we introduce another application of our main theorem. It is well-known that weakly special subvarieties associated to the period mapping $\D\Phi:B\rightarrow\Gamma\backslash$ D are algebraic subvarieties of $B$. The original version of this theorem is due to [CDK95] and in this form it was proven in [BKT20], [BBT22] using o-minimal geometry and the definability of period mappings, though it follows formally from [CDK95]. We will give yet another proof of this theorem using log-manifolds.
Regarding the necessary backgrounds, we refer readers to [KO21] or [Kli22] for a comprehensive overview.